# Blowouts of Nascent Wind Bubbles in Pulsar-Driven Supernovae

### v1

@@ II.3.1 Geometry and Scaling
$\rho(r)=\frac{M_{\rm adm}}{2\pi R_{\rm in}^{3}}\left(\frac{r}{R_{\rm in}}\right)^{-5}\left\{\left[(1-h)^{-2}-\left(\zeta_{\rm r}-h\right)^{-2}\right]\left[1-h\left(\frac{R_{\rm in}}{r}\right)\right]^{-3}\right\},$ (17)
@@ II.3.2 Evolution Equations
We next consider the evolution of the bulk kinetic energy. The kinetic energy of the admixture layer evolves through three channels: (i) work done at its boundaries, (ii) the conversion of internal energy into bulk motion through adiabatic pressure volume (pdVp\,dV) work, and (iii) advection of kinetic energy carried by upstream ejecta as it is entrained into the layer. At the inner boundary, the swept-up shell is accelerated by the overpressure of the pulsar-wind bubble, which applies an interface thermal pressure of order $p_{\rm in}$:
@@
At the outer boundary $R_{\rm out}$, because the admixture layer expands faster than the preshock ejecta, a forward shock forms and the associated ram pressure acts as a drag, hindering further expansion or thickening. However, hydrodynamical instabilities can fragment the interaction front into filamentary structures, so that strong shocks may occur only over part of the surface, while the intervening regions allow external material to enter with much weaker dissipation. As a result, the ram-pressure drag is effectively reduced. We therefore introduce a dimensionless shock covering fraction $\eta_{S}\!\in\![0,1]$ and write
@@
where $\rho_{\rm pre}\equiv\rho_{\rm ej}(R_{\rm out},t)$ is the preshock ejecta density given by the broken power-law profile, and
@@
is the velocity discontinuity at the shock (for a homologous preshock flow, $v_{\rm ej}(R_{\rm out},t)=R_{\rm out}/t$). The kinetic–energy loss associated with $F_{\rm out}$ defines the shock–heating power at $R_{\rm out}$,
@@
i.e., the rate at which bulk kinetic energy is extracted at $R_{\rm out}$ and converted into radiation, or into internal energy depending on the optical depth ahead.
@@
Assuming a radiation-dominated equation of state so that $p\propto\rho^{4/3}$ and adopting the affine velocity profile of Eq. (16), the integral evaluates to
$\dot{W}_{\rm pdV}\;=\;4\pi\,p_{\rm in}\,R_{\rm in}^{2}\,\frac{v_{\rm out}-v_{\rm in}}{\zeta_{\rm r}-1}\left[3\,\mathcal{G}\!\left(\zeta_{\rm r},h\right)-2h\,\mathcal{J}\!\left(\zeta_{\rm r},h\right)\right],$ (32)
where the function $\mathcal{J}$ is another Gauss hypergeometric function,
@@
In this convention, $\dot{W}_{\rm pdV}>0$ corresponds to a net conversion of internal energy into bulk kinetic energy during expansion.
In addition to the effects we discussed above, the bulk kinetic energy is also modified by the advection of kinetic energy carried by the upstream ejecta as it crosses $R_{\rm out}$ and is entrained into the layer. We write the corresponding kinetic luminosity as
$L_{\rm adv}\;\equiv\;\frac{1}{2}\,4\pi R_{\rm out}^{2}\,\rho_{\rm pre}\,\Delta v\,\bigl[v_{\rm ej}(R_{\rm out},t)\bigr]^{2}.$ (34)
This term counts the import of upstream kinetic energy irrespective of whether the material subsequently shocks or mixes quasi-continuously.
Collecting the terms above, the bulk kinetic energy evolves as
@@
which closes the kinetic–energy budget for the admixture layer under our sign conventions.
Turning to the momentum evolution, and retaining only the mechanical stresses at the boundaries while neglecting subdominant effects such as radiative momentum losses, we obtain
$\frac{dP_{\rm adm}}{dt}\;=\;F_{\rm in}-F_{\rm out}+4\pi\,R_{\rm out}^{2}\,\rho_{\rm ej}\,v_{\rm out}.$ (36)
We finally consider the internal energy budget before the admixture layer becomes optically thin. In addition to being converted into bulk motion through pdVp\,dV work, the internal energy is supplied by thermalization of the pulsar spin-down power and depleted by photon leakage at $R_{\rm out}$. Denoting by $\epsilon_{\rm adm}$ the fraction of $L_{\rm sp}(t)$ that is thermalized within the layer, we write
@@
$\epsilon_{\rm adm}=1-\exp(-\tau_{\rm ab}),\qquad\tau_{\rm ab}=\int_{R_{\rm in}}^{R_{\rm out}}\kappa_{\rm ab}\,\rho\,dr.$ (38)
Before the admixture layer becomes transparent $\epsilon_{\rm adm}\simeq 1$. In this regime we evaluate $\tau_{\rm ab}$ using a simplified effective absorption opacity $\kappa_{\rm ab}$, which we treat as an approximate measure of the optical depth for absorbed radiation. The $(1-f_{\rm esc})L_{\rm sh}$ term accounts for the fraction of the shock-heating luminosity that is trapped within the admixture layer when photon transport in the exterior is inefficient, so that photons produced at the forward shock are overtaken by the advancing boundary. Technically, this term is $\simeq 0$, in most cases of our calculation. See Eq. (39) and the accompanying discussion. One may notice that the energy-injection term in Eq. (37) does not include contributions from the radioactive decay of 56Ni. In the blowout-enabled regime considered here, the pulsar spin-down power is expected to dominate the energy budget and to set the dynamics of the admixture layer. We therefore neglect radioactive heating in the present work to simplify the model and to focus on the dominant pulsar-driven effects. Incorporating 56Ni decay and its coupling to the layer will be deferred to a more comprehensive follow-up model.
While the diffusion front is still beyond $R_{\rm out}$, the local expansion speed exceeds the effective diffusion speed throughout the layer, so photon transport is predominantly advective and outward leakage is inefficient. Nevertheless, a finite leakage exists and grows as the diffusion front approaches $R_{\rm out}$. To capture this, we model the escape at $R_{\rm out}$ as a competition between diffusion and boundary advection, characterized by an effective diffusion speed $v_{\rm diff}$ and an outward boundary speed $v_{\rm out}$. The probability that a photon effectively outruns the moving boundary is then parameterized as
@@
where the factor of 2 accounts for the forward–hemisphere excess (backward scattering being subdominant in the outward-biased radiation field). With this prescription, the diffusive luminosity through $R_{\rm out}$ is
@@
This construction reproduces the standard diffusion–luminosity scaling $L_{\rm diff}\propto R_{\rm out}^{2}\,[c/(\kappa\rho(R_{\rm out}))]\,|\partial_{r}p|$, while allowing a smooth transition from the advection–dominated to the leakage–dominated regime as the diffusion front reaches $R_{\rm out}$.
An analogous consideration applies to the radiation produced at the forward shock. Before diffusion front penetrates into the admixture layer, the forward shock is located inside the photosphere ($\tau\!\sim\!1$) surface so that the upstream is optically thick ($\tau\gtrsim 1$), the shock is radiation–mediated. However, a non-negligible fraction of the shock-generated photons does not escape immediately; instead, they are captured by the following optically thick, fast expanding blowout admixture layer frontier. Consistent with this, we assign a fraction $(1-f_{\rm esc})$ of the shock–heating power $L_{\rm sh}$ to the internal-energy reservoir, as represented by the final term in Eq. (37).
@@ II.3.3 Admixture Layer Transparency
Motivated by the contrast in optical depth and thermodynamic state across the advancing diffusion front, we treat the layer as an optically thick interior and an optically thin exterior separated by $R_{\rm diff}$. Radiative leakage ahead of the front weakens pressure support on a dynamical timescale, whereas the interior remains in near LTE and is well described by an approximately adiabatic $\gamma\simeq 4/3$ equation of state. (i) The optically thick inner zone, $R_{\rm in}\leq r\leq R_{\rm diff}$, is described by the same structural profiles and equation of state as in the previous subsection, with the replacement $R_{\rm out}\!\rightarrow\!R_{\rm diff}$ (equivalently, $\zeta_{\rm r}\!\rightarrow\!\zeta_{\rm d}\!\equiv\!R_{\rm diff}/R_{\rm in}$). (ii) The optically thin outer zone, $R_{\rm diff}\leq r\leq R_{\rm out}$, cools rapidly by diffusion and exerts only a subdominant dynamical back–reaction prior to global transparency. Once $R_{\rm diff}$ has penetrated into the admixture layer, a strong forward shock at $R_{\rm out}$ is no longer sustained. The structure between $R_{\rm diff}$ and $R_{\rm out}$ therefore should not rapidly deviate from that at the moment the diffusion front passes. In the absence of strong external forcing capable of rapidly reshaping the stratification, we thus treat the optically thin exterior as kinematically frozen at that time. For density-dependent calculations, such as optical-depth estimates, we approximate $R_{\rm diff}\leq r\leq R_{\rm out}$ by extending the instantaneous thick-region density profile beyond $R_{\rm diff}$, retaining $\rho\propto r^{-5}$.
With this partition, we write the state evolution of the optically thin zone. Let $\epsilon_{\rm th,\,thick}$ and $\epsilon_{\rm th,\,thin}$ denote the thermalization efficiencies of the pulsar spin–down power $L_{\rm sp}(t)$ within the thick and thin zones, respectively. The thin–zone internal energy evolves as
@@
$L_{\rm th}\;=\;\frac{E_{\rm int,\,thin}}{t_{\rm esc,\,thin}},$ (46)
@@ II.3.4 Observable Signatures
Having specified the evolution from blowout to global transparency, we now connect the model to observables. We have distinguished three contributing channels: (i) shock heating at the forward shock, (ii) photon diffusion from the admixture layer, and (iii) graybody emission from the thermalized post-diffusion cold region. Crucially, especially for the shock–heating and photon diffusion components our expressions quantify local photon losses at the emission site; they do not yet include radiative transfer through the overlying ejecta to the observer. During the intervals when these channels dominate, the photosphere $R_{\rm ph}$ (where $\tau\!\sim\!1$) typically lies exterior to the local radiating surface. The region beyond emission sites, although may relatively optically thin, still possess a finite optical depth. Photons that leave the shock region or decouple at the diffusion front can therefore undergo additional scatterings in this residual scattering envelope.
To capture the time delay introduced by residual scatterings between the emission site and the observer, we introduce a “photon pool” spanning the region between a local emission radius $R_{\rm emit}$ and the last-scattering surface $R_{\rm ph}$. This prescription is intended to model the scattering-induced hold-up after photons are released locally but before they free-stream from the photosphere. In our setup, the relevant local sources are the forward-shock heating luminosity $L_{\rm sh}$ (emitted at $R_{\rm out}$) and the diffusive leakage from the hot admixture layer $L_{\rm diff}$ (emitted at $R_{\rm diff}$). In Eq. (49) we denote either contribution generically as an injected power $L_{\rm inj}$, but we evolve separate photon pools for these two components because their emission radii differ and therefore their scattering delays cannot, in general, be combined into a single effective pool. By contrast, emission from the post-diffusion cold region is released once the exterior is already optically thin, so an additional scattering delay is negligible and we do not apply the pool treatment there. Photons that cross $R_{\rm ph}$ are counted as escaped to infinity.
@@
The transient storage of radiation within $R_{\rm emit}\!\lesssim\!r\!\lesssim\!R_{\rm ph}$ is represented by a pool energy $E_{\rm pool}(t)$, which evolves as
$\frac{dE_{\rm pool}}{dt}\;=\;L_{\rm inj}\;-\;L_{\rm scat},$ (49)
where $L_{\rm inj}$ is the local radiative power injected into the pool and
$L_{\rm scat}\;=\;\frac{E_{\rm pool}}{t_{\rm scat}}$ (50)
@@
is the emergent luminosity after scatterings. Here $t_{\rm scat}$ is the characteristic scattering (escape) timescale across the layer between $R_{\rm emit}$ and $R_{\rm ph}$. Because the pool region is effectively optically thin and contains little inertia, pdVp\,dV work is negligible; its primary role is to delay and temporally broaden the emergent signal. Since the shock- and diffusion-powered emission sites are generally not co-spatial, we treat their photon pools independently.
@@
where the subscript ‘scat’ denotes the luminosity after passage through the scattering envelope. The corresponding effective blackbody temperature is

### v2

@@ II.2 Blowout Condition
The discussion above presumes that the swept-up shell remains quasi-adiabatic and geometrically thin until the forward shock reaches $R_{\rm tr}$ (Chevalier and Fransson, 1992). Equivalently, the diffusion front, the radius where radiation and matter begin to decouple and diffusion becomes dynamically important, must not overtake the shell before it reaches $R_{\rm tr}$. In practice, this is captured by requiring $t_{\rm bo}<t_{\rm diff}$ (Eqs. (10)–(11)); otherwise the system transitions to an inner-breakout–like evolution (Kasen et al., 2016) and the thin-shell approximation breaks down.
Across the parameter space enclosed by Eqs. (8) and (11) in Fig. 2, this additional consistency condition is satisfied except near the luminosity threshold $L_{\rm sp}\sim E_{\rm bo}/t_{\rm diff}$, where $t_{\rm bo}\sim t_{\rm diff}$ and diffusion becomes marginally important already during the pre-blowout phase. This corresponds to the marginal-blowout regime discussed in Section IV.2.
@@ II.3.1 Geometry and Scaling
$\rho(r)=\frac{M_{\rm adm}}{2\pi R_{\rm in}^{3}}\left(\frac{r}{R_{\rm in}}\right)^{-5}\left\{\left[(1-h)^{-2}-\left(\zeta_{\rm r}-h\right)^{-2}\right]^{-1}\left[1-h\left(\frac{R_{\rm in}}{r}\right)\right]^{-3}\right\},$ (17)
@@ II.3.2 Evolution Equations
We evolve the bulk kinetic energy of the admixture layer by accounting for (i) boundary work, (ii) conversion between internal and kinetic energy via adiabatic pdVp\,dV work, and (iii) kinetic-energy advection by upstream ejecta entrained through the outer boundary.
At the inner boundary, the pulsar-wind bubble overpressurizes the layer and exerts an outward thrust set by the interface thermal pressure $p_{\rm in}$,
@@
At the outer boundary $R_{\rm out}$, the layer outruns the preshock homologous ejecta and drives a forward shock. The associated ram pressure acts as a drag, but post-blowout fragmentation can render the interaction front porous so that strong shocks cover only a fraction of the surface. We parameterize this reduction by a dimensionless shock covering fraction $\eta_{S}\!\in\![0,1]$ and write
@@
where $\rho_{\rm pre}\equiv\rho_{\rm ej}(R_{\rm out},t)$ is the preshock ejecta density from the broken power-law profile, and
@@
is the velocity jump at the shock (for a homologous preshock flow, $v_{\rm ej}(R_{\rm out},t)=R_{\rm out}/t$). The kinetic-energy extraction at $R_{\rm out}$ defines the shock-heating power,
@@
i.e., the rate at which bulk kinetic energy is converted into heat (and ultimately radiation, depending on the optical depth ahead).
@@
Assuming a radiation-dominated equation of state, $p\propto\rho^{4/3}$, and adopting the affine velocity profile of Eq. (16), the integral evaluates to
$\dot{W}_{\rm pdV}\;=\;4\pi\,p_{\rm in}\,R_{\rm in}^{2}\,\frac{v_{\rm out}-v_{\rm in}}{\zeta_{\rm r}-1}\left[3\,\mathcal{G}\!\left(\zeta_{\rm r},h\right)-2h\,\mathcal{J}(\zeta_{\rm r},h)\right],$ (32)
where $\mathcal{J}$ is another Gauss hypergeometric function,
@@
With this convention, $\dot{W}_{\rm pdV}>0$ corresponds to a net transfer of internal energy into bulk kinetic energy during expansion.
Mass loading also imports upstream kinetic energy into the layer as preshock ejecta crosses $R_{\rm out}$. We write the associated kinetic luminosity as
$L_{\rm adv}\;\equiv\;\frac{1}{2}\,4\pi R_{\rm out}^{2}\,\rho_{\rm pre}\,\Delta v\,\bigl[v_{\rm ej}(R_{\rm out},t)\bigr]^{2},$ (34)
which counts the inflow of upstream kinetic energy irrespective of whether the entrained material subsequently shocks strongly or mixes quasi-continuously.
Collecting the above contributions, the bulk kinetic energy evolves as
@@
closing the kinetic-energy budget for the admixture layer under our sign conventions.
Turning to the momentum evolution, and retaining only mechanical stresses at the boundaries (neglecting subdominant radiative momentum losses), we obtain
$\frac{dP_{\rm adm}}{dt}\;=\;F_{\rm in}-F_{\rm out}+4\pi\,R_{\rm out}^{2}\,\rho_{\rm pre}\,\Delta v\,v_{\rm out}.$ (36)
In Eq. (36), $F_{\rm in}-F_{\rm out}$ is the net boundary thrust driving bulk acceleration. The last term is the post-shock momentum flux associated with mass loading: ejecta is swept into the layer at the rate $\dot{M}=4\pi R_{\rm out}^{2}\rho_{\rm pre}\Delta v$ and is rapidly brought to co-motion with the outer boundary, so the momentum added to the layer is $\dot{M}\,v_{\rm out}$.
We finally consider the internal-energy budget while the admixture layer remains optically thick. In addition to pdVp\,dV conversion, internal energy is supplied by thermalization of the pulsar spin-down power and is depleted by radiative leakage through $R_{\rm out}$. Denoting by $\epsilon_{\rm adm}$ the fraction of $L_{\rm sp}(t)$ thermalized within the layer, we write
@@
$\epsilon_{\rm adm}=1-\exp(-\tau_{\rm ab}),\qquad\tau_{\rm ab}=\int_{R_{\rm in}}^{R_{\rm out}}\kappa_{\rm ab}\,\rho\,dr,$ (38)
which approaches unity while the layer is optically thick. In this regime we evaluate $\tau_{\rm ab}$ using a simplified effective absorption opacity $\kappa_{\rm ab}$, treated as a proxy for the depth over which injected power is absorbed and thermalized. The $(1-f_{\rm esc})L_{\rm sh}$ term accounts for the fraction of shock-generated radiation that is trapped because photon transport in the exterior is inefficient: photons produced near $R_{\rm out}$ can be overtaken by the advancing boundary and reprocessed within the layer. Technically, this term is $\simeq 0$ in most cases of our calculation; see Eq. (39) and the accompanying discussion.
Eq. (37) neglects energy injection from 56Ni decay for simplicity. In the blowout branch considered here, the engine injects an energy comparable to the explosion energy and well above the radioactive contribution within short time, which justifies ignoring 56Ni heating in most cases. While, we note that 56Ni decay can still produce a visible effect in the marginal blowout regime and to the post main peak observable signatures. We summarize these effects in Section IV.3 and will incorporate a self consistent treatment of 56Ni decay in future extensions of the model.
While the diffusion front remains beyond $R_{\rm out}$, expansion outpaces diffusion across the layer, so photon transport is predominantly advective and direct leakage is weak. To capture the gradual onset of leakage as the diffusion front approaches $R_{\rm out}$, we model escape through $R_{\rm out}$ as a competition between diffusion and boundary advection, characterized by an effective diffusion speed $v_{\rm diff}$ and the outward boundary speed $v_{\rm out}$. The probability that a photon outruns the moving boundary is parameterized as
@@
where the factor of 2 approximates the forward-hemisphere excess in an outward-biased radiation field. The diffusive luminosity through $R_{\rm out}$ is then
@@
This construction reproduces the standard scaling $L_{\rm diff}\propto R_{\rm out}^{2}\,[c/(\kappa\rho(R_{\rm out}))]\,|\partial_{r}p|$, while providing a smooth transition from the advection-dominated regime to the leakage-dominated regime as the diffusion front reaches $R_{\rm out}$.
An analogous consideration applies to radiation produced at the forward shock. Before the diffusion front penetrates into the admixture layer, the forward shock lies inside the photosphere ($\tau\!\sim\!1$) so that the upstream remains optically thick ($\tau\gtrsim 1$) and the shock is radiation-mediated. A non-negligible fraction of the shock-generated photons is therefore captured by the following optically thick, rapidly expanding blowout frontier; accordingly, we deposit a fraction $(1-f_{\rm esc})$ of $L_{\rm sh}$ into the internal-energy reservoir, as represented by the final term in Eq. (37).
@@ II.3.3 Admixture Layer Transparency
Motivated by the contrast in optical depth and thermodynamic state across the advancing diffusion front, we treat the layer as an optically thick interior and an optically thin exterior separated by $R_{\rm diff}$. Radiative leakage ahead of the front weakens pressure support on a dynamical timescale, whereas the interior remains in near LTE and is well described by an approximately adiabatic $\gamma\simeq 4/3$ equation of state. (i) The optically thick inner zone, $R_{\rm in}\leq r\leq R_{\rm diff}$, is described by the same structural profiles and equation of state as in the previous subsection, with the replacement $R_{\rm out}\!\rightarrow\!R_{\rm diff}$ (equivalently, $\zeta_{\rm r}\!\rightarrow\!\zeta_{\rm d}\!\equiv\!R_{\rm diff}/R_{\rm in}$). (ii) The optically thin outer zone, $R_{\rm diff}\leq r\leq R_{\rm out}$, cools rapidly by diffusion and exerts only a subdominant dynamical back–reaction prior to global transparency.
Once $R_{\rm diff}$ has penetrated into the admixture layer, a strong forward shock at $R_{\rm out}$ is no longer sustained. The structure between $R_{\rm diff}$ and $R_{\rm out}$ therefore should not rapidly deviate from that at the moment the diffusion front passes. In the absence of strong external forcing capable of rapidly reshaping the stratification, we thus treat the optically thin exterior as kinematically frozen at that time. For density-dependent calculations, such as optical-depth estimates, we approximate $R_{\rm diff}\leq r\leq R_{\rm out}$ by extending the instantaneous thick-region density profile beyond $R_{\rm diff}$, retaining $\rho\propto r^{-5}$.
With this partition, we write the state evolution of the optically thin zone, using a one-zone approximation, since pressure and temperature gradients are dynamically unimportant once photons decouple. Let $\epsilon_{\rm th,\,thick}$ and $\epsilon_{\rm th,\,thin}$ denote the thermalization efficiencies of the pulsar spin–down power $L_{\rm sp}(t)$ within the thick and thin zones, respectively. The thin–zone internal energy evolves as
@@
$L_{\rm thin}\;=\;\frac{E_{\rm int,\,thin}}{t_{\rm esc,\,thin}},$ (46)
@@ II.3.4 Observable Signatures
Having specified the evolution from blowout to global transparency, we now connect the model to observables. We have distinguished three contributing channels: (i) shock heating at the forward shock, (ii) photon diffusion from the admixture layer, and (iii) graybody emission from the thermalized post-diffusion cold region. Crucially, especially for the shock–heating and photon diffusion components our expressions quantify local photon losses at the emission site; they do not yet include radiative transfer through the overlying ejecta to the observer. During the intervals when these channels dominate, the photosphere $R_{\rm ph}$ (where $\tau\!\sim\!1$) typically lies exterior to the local radiating surface. The region beyond emission sites, although it may be relatively optically thin, still possesses a finite optical depth. Photons that leave the shock region or decouple at the diffusion front can therefore undergo additional scatterings in this residual scattering envelope.
In the residual-scattering envelope, photons have already largely decoupled from the gas and the transport is near the free-streaming limit ($\tau\!\lesssim\!{\rm few}$). The finite optical depth between a local emission radius $R_{\rm emit}$ and the last-scattering surface $R_{\rm ph}$ therefore acts primarily as a modest scattering-induced hold-up, delaying and temporally smoothing photons after they are released locally but before they free-stream from the photosphere. We model this effect with a one-zone residual-scattering reservoir, introduced solely to capture this small propagation delay, while leaving the stratification-dependent diffusion that sets $R_{\rm diff}$ and $L_{\rm diff}$ inside the optically thick ejecta unchanged.
In our setup the relevant local sources are the forward-shock heating luminosity $L_{\rm sh}$ (emitted at $R_{\rm out}$) and the diffusive leakage from the hot admixture layer $L_{\rm diff}$ (emitted at $R_{\rm diff}$); we evolve separate reservoirs because their emission radii differ and their scattering delays cannot, in general, be collapsed into a single effective one. By contrast, emission from the post-diffusion cold region is released effectively at $R_{\rm ph}$, so there is no overlying scattering envelope to traverse; any additional scattering delay is therefore negligible and we do not apply this treatment there. Importantly, near the first-peak maximum the outer layers are already tenuous and $R_{\rm ph}\simeq R_{\rm out}$, so the residual-scattering layer is geometrically thin and of low inertia; the one-zone closure is therefore sufficient and does not distort the key first-peak signature, nor does it bias the timing or luminosity of the subsequent main peak.
@@
The transient storage of radiation within $R_{\rm emit}\!\lesssim\!r\!\lesssim\!R_{\rm ph}$ is represented by a reservoir energy $E_{\rm rs}(t)$, which evolves as
$\frac{dE_{\rm rs}}{dt}\;=\;L_{\rm inj}\;-\;L_{\rm scat},$ (49)
where $L_{\rm inj}$ is the local radiative power injected into the reservoir and
$L_{\rm scat}\;=\;\frac{E_{\rm rs}}{t_{\rm scat}}$ (50)
@@
is the emergent luminosity after residual scatterings. Here $t_{\rm scat}$ is the characteristic scattering (escape) timescale across the layer between $R_{\rm emit}$ and $R_{\rm ph}$. Because the residual-scattering region contains little inertia, pdVp\,dV work is negligible; its primary role is to delay and temporally broaden the emergent signal. Since the shock- and diffusion-powered emission sites are generally not co-spatial, we treat their residual-scattering reservoirs independently.
@@
where the subscript ‘scat’ denotes the luminosity after passage through the residual scattering envelope. The corresponding effective blackbody temperature is
@@
With estimates for the admixture-layer outer radius at the first peak (Eq. (55)) and the corresponding bolometric luminosity (Eqs. (63)–(64)), we can obtain a compact prediction for the characteristic color temperature of the first peak. By construction, the first maximum occurs when the diffusion front has just entered the admixture layer, i.e. when $R_{\rm diff}$ approaches the outer edge. At this epoch the ejecta exterior to the diffusion front is already close to optically thin, so the key radii are well approximated by $R_{\rm diff}\sim R_{\rm ph}\sim R_{\rm out}$. We therefore estimate the first-peak effective temperature using $R_{\rm out}(t_{\rm first})$ as the photospheric scale,
$T_{\rm eff,1}\;\equiv\;\left(\frac{L_{\rm first}}{4\pi\sigma R_{\rm out}^{2}(t_{\rm first})}\right)^{1/4}\simeq 1.61\times 10^{5}\ {\rm K}\;\left(\frac{E_{\rm kin,adm}}{10^{51}\ {\rm erg}}\right)^{3/16}\left(\frac{M_{\rm ej}}{10\,M_{\odot}}\right)^{-5/16}\left(\frac{\kappa}{0.1\ {\rm cm^{2}\,g^{-1}}}\right)^{-3/8},$ (65)
Eq. (65) implies $T_{\rm eff,1}\sim 10^{5}\,$K for fiducial SESNe parameters, thus, the first peak should therefore be predominantly UV, and any contemporaneous optical emission is expected to be extremely blue. In the most extreme cases the optical bands probe the Rayleigh–Jeans tail of the spectrum, naturally yielding a comparatively muted or even visually inconspicuous first peak in optical light curves despite a large bolometric output.